OpenStudy (anonymous):

A computer is programmed to generate a sequence of three digits, where each digit is either 0 or 1, and each of these is equally likely to occur. Construct a sample space that shows all possible three-digit sequences of 0s and 1s and then find the probability that a sequence will contain exactly one 0.

OpenStudy (anonymous):

total outcome = 2^3 = 8 exactly one 0 = 3C1 = 3 so 3/8

jimthompson5910 (jim_thompson5910):

The sample space is the set of all possible outcomes. So in this case, the sample space is 000 001 010 100 011 101 110 111 Notice how there are 3 instances where you have exactly 1 zero (011,101, 110) out of 8 total possible outcomes. So that's why the answer is 3/8

OpenStudy (anonymous):

okay thank u so much!!!!!!!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

The manager of a convenience store wishes to determine how many cartons of eggs are damaged in shipment and delivery. For ten shipments of 1000 cartons of eggs, she examines every 50th carton to see how many cartons contain cracked eggs. How large is the sample?

OpenStudy (anonymous):

a: 200 cartons b: the 50th carton c:1000 cartons d: cannot determine from this information

jimthompson5910 (jim_thompson5910):

oh sorry, i didn't see the "ten" part there are 10*1000 = 10,000 cartons total so the sample size is 10,000/50 = 200 cartons

OpenStudy (anonymous):

There are three schedule options for classes at a high school: 90-minute classes every other day for a year, 90-minute classes every day for a semester, or 45-minute classes every day for a year. Out of 1200 students, 50 students from each grade level are chosen at random and asked their preference. Identify the population and the sample.

jimthompson5910 (jim_thompson5910):

were you able to get anywhere?

OpenStudy (anonymous):

yes it was correct. your being a life savor because i am the worst at math and i need to get these right for my online e2020 class

jimthompson5910 (jim_thompson5910):

I meant for the current problem. Were you able to get anywhere with it?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

the info will be drawn from this sentence "Out of 1200 students, 50 students from each grade level are chosen at random and asked their preference"

jimthompson5910 (jim_thompson5910):

the population is ALWAYS larger than the sample size

jimthompson5910 (jim_thompson5910):

so the population is the 1200 students the sample size is the 50 selected

jimthompson5910 (jim_thompson5910):

"50 students from each grade level " I mean, so if you're talking about 4 grade levels, then you're talking about 4*50 = 200 students

OpenStudy (anonymous):

oh ok. i understand that.

OpenStudy (anonymous):

Is the data set “color of the coal taken from a railcar” quantitative or qualitative? If it is quantitative, is it discrete or continuous?

jimthompson5910 (jim_thompson5910):

is color a number in any way?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

so it has to be qualitative

jimthompson5910 (jim_thompson5910):

if you can put it in some numerical way (and do math on it), then it's quantitative (think quantity)

OpenStudy (anonymous):

ahhh ok i never knew what those were till now.

OpenStudy (anonymous):

what about discrete r continous

jimthompson5910 (jim_thompson5910):

notice how it says "If it is quantitative, is it discrete or continuous? " IF it's quantitative...it's not...so we don't have to worry about discrete vs continuous

OpenStudy (anonymous):

i know that but what are they? when do you know which is which?

jimthompson5910 (jim_thompson5910):

discrete data is something you can count (example: the number of people in the classroom) continuous data is something that's measured, weighed, etc (example: heights of students) with discrete data, there are gaps in the numbers and the numbers are usually whole numbers (ex: there's a gap between 1 student and 2 students) with continuous data, there are no gaps in the numbers (you can be between 1 and 2 to get say 1.5)

OpenStudy (anonymous):

calories burned while running a marathon?

jimthompson5910 (jim_thompson5910):

you can burn 1 calorie, 1/2 a calorie, 0.2 calories, etc etc so this is continuous data

OpenStudy (anonymous):

distance a baseball travels after being hit would be continuous?

jimthompson5910 (jim_thompson5910):

yep since you can have 20 feet, 20.2 ft, 20.0002 ft, etc

OpenStudy (anonymous):

what about the number of leaves on a branch? that would be counting so its discrete?

jimthompson5910 (jim_thompson5910):

you nailed it

jimthompson5910 (jim_thompson5910):

you can't have half a leaf (well technically you can,but we're counting the number of WHOLE leaves) or anything like that

OpenStudy (anonymous):

haha ok i get the hang of that now. those i thought were hard but i now see are so simple.

jimthompson5910 (jim_thompson5910):

yeah they're not too bad